Session: Sep 1997 (session 3)
Frequency: 3.6/13 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Cattani at Sat Sep 27 20:32:31 METDST 1997) Success

Onsala (Antonis Polatidis at Sat Sep 27 23:56:44 METDST 1997) Success
We had problems with the positioning of the Head 2 (read head). This should not affect the quality of the data. Otherwise no known problems. Very nice weather.

Effelsberg (dgraham at Mon Sep 29 11:55:30 METDST 1997) Minor failures
Lost 0800-0935 due to computer problem.

Noto (C.Trigilio at Mon Sep 29 14:24:07 METDST 1997) Success

OAN\-Yebes (P. de Vicente at Mon Sep 29 16:03:21 METDST 1997) Minor failures
Continous formatter problems. MKIII sampler used

Westerbork (Tony Foley at Tue Sep 30 16:03:54 METDST 1997) Minor failures
Lots of false alarms abou tape movement and some sources out of HA range. System had only 1 linear polarization (my apologies but 1 of the channels was dead, so we could not make circular) While rewinding at the end the capstan lost grip and the tape was slipping

Calibration (Fredrik T Rantakyro at Wed Nov 5 16:05:02 MET 1997) Minor failures
The ANTAB formatted TSYS information can be found on VLBEER in the SEP97 directory. The file is called gg032b.antab and contains the latest antenna calibration information.